### v2

# Kang Chen♣,♡,∗, Zhihao Liu♢,♡,∗, Tonghe Zhang♯,⊳,∗, Zhen Guo, Si Xu, Hao Lin, Hongzhi Zang, & Xiang Li, Quanlu Zhang, Zhaofei Yu, Guoliang Fan, Tiejun Huang, Yu Wang♠,†, Chao Yu♠,♡,†

Abstract
Vision-Language-Action (VLA) models enable robots to understand and perform complex tasks from multimodal input. Although recent work explores using reinforcement learning (RL) to automate the laborious data collection process in scaling supervised fine-tuning (SFT), applying large-scale RL to flow-based VLAs (e.g., $\pi_{0}$, $\pi_{0.5}$) remains challenging due to intractable action log-likelihoods from iterative denoising. We address this challenge with $\pi_{\texttt{RL}}$, an open-source framework for training flow-based VLAs in parallel simulation. $\pi_{\texttt{RL}}$ implements two RL algorithms: (1) Flow-Noise models the denoising process as a discrete-time MDP with a learnable noise network for exact log-likelihood computation. (2) Flow-SDE integrates denoising with agent-environment interaction, formulating a two-layer MDP that employs ODE-to-SDE conversion for efficient RL exploration. We evaluate $\pi_{\texttt{RL}}$ on LIBERO, ManiSkill, and MetaWorld benchmarks. On LIBERO, $\pi_{\texttt{RL}}$ boosts few-shot SFT models $\pi_{0}$ and $\pi_{0.5}$ from 57.6% to 97.6% and from 77.1% to 98.3%, respectively. On ManiSkill, we train $\pi_{\texttt{RL}}$ in 320 parallel environments, improving $\pi_{0}$ from 38.4% to 78.8% and $\pi_{0.5}$ from 40.1% to 90.8% across 4352 variations of pick-and-place task. On MetaWorld, RL is conducted over 50 different manipulation tasks and yields performance gains of 35.0% and 26.9% for $\pi_{0}$ and $\pi_{0.5}$ models, respectively. Overall, $\pi_{\texttt{RL}}$ achieves significant performance gains and stronger generalization over SFT-models, validating the effectiveness of online RL for flow-based VLAs.
Figure 1: Overview of $\pi_{\texttt{RL}}$. $\pi_{\texttt{RL}}$, an online RL framework featuring Flow-Noise and Flow-SDE two approaches, is designed to enhance the performance and generalization of SFT-aligned flow-based VLAs, represented by the $\pi_{0}$ and $\pi_{0.5}$. Experiments conducted on LIBERO, ManiSkill, and MetaWorld benchmarks demonstrate that $\pi_{\texttt{RL}}$ achieves significant gains over SFT models.
1 Introduction
Vision-Language-Action (VLA) models (din2025vision) have emerged as a leading solution for general-purpose robots, effectively bridging the gap between high-level multimodal reasoning and low-level physical control (firoozi2025foundation). Conditioned on sensor inputs and language commands, VLAs (team2024octo; kim2024openvla; black2024pi_0; intelligence2025pi05) can translate abstract instructions into executable robotic actions, thereby enabling intuitive and flexible human-robot interaction.
The training methodology for VLAs follows the standard pre-training and supervised fine-tuning (SFT) paradigm as shown in Fig.˜1. Building on the pretrained Vision-Language Model (VLM) (touvron2023llama; beyer2024paligemma), VLAs are fine-tuned on large-scale, heterogeneous human demonstration datasets (o2024open; khazatsky2024droid), followed by SFT on the target task to align their capabilities with the specific embodiment and environment. However, reliance on SFT introduces a critical challenge: curating large-scale, high-quality expert trajectories is both laborious and costly (din2025vision), and the models obtained via SFT tend to overfit to expert demonstrations (liberoplus).
Recent efforts (zang2025rlinf; li2025simplevla; tan2025riptvla; rl4vla) have explored expanding the VLA training process with reinforcement learning (RL), establishing a pre-training, SFT, and RL paradigm as shown in Fig.˜1, allowing VLAs to improve their performance beyond initial expert demonstrations through active environmental interaction and the development of more generalizable policies.
However, these RL advances have been largely confined to autoregressive VLAs, featuring OpenVLA (kim2024openvla) and OpenVLA-OFT (kim2025openvlaoft), which employ discrete action decoders that generate output in an autoregressive or parallel fashion. This stands in stark contrast to diffusion- or flow-based VLAs, exemplified by the $\pi$ series models $\pi_{0}$ (black2024pi_0) and $\pi_{0.5}$ (intelligence2025pi05), which generate actions through iterative refinement in flow matching (lipman2022flow), offering the advantages of generating action chunks in high-frequency and performing highly dexterous tasks (black2024pi_0). Consequently, previous VLA-RL algorithms are incompatible with flow-based VLAs, and the fundamental challenge lies in how to characterize a logarithmic likelihood (hutchinson1989stochastic; chen2018neural) for the executed actions.
Contribution: We introduce $\pi_{\texttt{RL}}$, the first open-source framework for fine-tuning flow-based VLAs $\pi_{0}$ and $\pi_{0.5}$ with parallel online RL.
To address the intractable log-likelihood estimation problem in flow matching, we propose two solutions: Flow-Noise and Flow-SDE. Flow-Noise integrates a learnable noise network into the denoising process and models this stage as a discrete-time Markov decision process (MDP) for exact log-likelihood estimation. Flow-SDE converts the ordinary differential equation (ODE) denoising process into a stochastic differential equation (SDE) while maintaining equivalent marginal distributions for exploration, and builds a two-layer MDP that couples the denoising process with policy-environment interaction, along with a hybrid ODE-SDE sampling technique for training acceleration. Given the formulated MDP and the exact log-likelihood computation, $\pi_{\texttt{RL}}$ undergoes further optimization via the proximal policy optimization (PPO) (schulman2017ppo) algorithm.
We conduct extensive experiments on the challenging multi-task benchmarks LIBERO (liu2023libero) and high-fidelity simulator ManiSkill (tao2024maniskill3) to evaluate the effectiveness of $\pi_{\texttt{RL}}$ optimization on the $\pi_{0}$ and $\pi_{0.5}$ models, with comprehensive findings summarized in Fig.˜1.
Results on LIBERO. $\pi_{\texttt{RL}}$ demonstrates substantial performance gains over the SFT baselines, with the average success rate of $\pi_{0}$ improving from 57.6% to 97.6%, and $\pi_{0.5}$ from 77.1% to 98.3%. Notably, on the LIBERO-Long task suite, $\pi_{\texttt{RL}}$ boosts the performance of the $\pi_{0.5}$ one-trajectory SFT model from 43.9% to 94.0%, surpassing the 92.4% performance of the all-trajectories SFT model. Moreover, we compare against group relative policy optimization (GRPO) (shao2024grpo) as an alternative policy gradient algorithm, with the comparison showing that PPO consistently outperforms GRPO across all task suites.
Results in ManiSkill. We train the policy to pick 16 types of objects and place them on 17 different receptacles in 16 photorealistic scenes, with a total of 4,352 combinations. $\pi_{\texttt{RL}}$ boosts the average success rate from 41.6% to 85.7% for $\pi_{0}$ and 41.9% to 84.8% for $\pi_{0.5}$, demonstrating $\pi_{\texttt{RL}}$’s ability to support large-scale multi-task RL. Additionally, we also conduct experiments on the SIMPLER benchmark (SIMPLER), the success rate was elevated from 67.2% to 86.7% for $\pi_{0}$ and from 59.24% to 79.1% for $\pi_{0.5}$.
Results in MetaWorld. Beyond the pick-and-place tasks detailed previously, we evaluate $\pi_{\texttt{RL}}$ on the MetaWorld MT50 benchmark to assess the capabilities of $\pi_{\texttt{RL}}$ across 50 diverse manipulation tasks of varying difficulty. With RL, the $\pi_{0}$ and $\pi_{0.5}$ models achieve success rates of 85.8% and 70.7%, exceeding the performance of the leading baseline SmolVLA (68.2%) (shukor2025smolvla).
To sum up, our contributions are:
  • RL for flow-based VLAs. We introduce $\pi_{\texttt{RL}}$, the first online RL fine-tuning framework for flow-based $\pi$-series VLAs, featuring Flow-Noise and Flow-SDE, two distinct technical solutions that allow exact log-likelihood estimation in flow matching.
  • Superior Performance. We demonstrate significant performance improvements and enhanced generalization of $\pi_{\texttt{RL}}$ on the multi-task benchmarks LIBERO and ManiSkill.
  • Comprehensive Ablation. We conduct thorough ablation studies on RL algorithms, critic designs, noise injection strategies, MDP formulations, and hyperparameters within flow-based VLAs, providing empirical insights for future research on RL for flow-based VLAs.
  • Open-source Code and Models. We release all codes and model checkpoints to ensure reproducibility, hope thating that our study helps to advance further research in this field.
2 Related Work
2.1 Vision-Language-Action Models
VLA models have recently achieved remarkable progress in robotics by integrating multimodal inputs to enable unified perception, reasoning, and control. This development has led to a series of architectures, including Octo (team2024octo), RT (brohan2022rt), OpenVLA, OpenVLA-OFT, $\pi_{0}$, $\pi_{0.5}$, and GR00T (bjorck2025gr00t). OpenVLA, which exemplifies the autoregressive VLA architecture, discretizes the action space into tokenized representations. This enables language-conditioned control by treating actions as part of the VLM’s vocabulary, but it inherently limits the resolution required for fine-grained motion. To achieve more dexterous and continuous physical behaviors, $\pi_{0}$ and $\pi_{0.5}$, as representatives of flow-based VLA architectures, introduce an action chunking architecture based on flow matching. This allows VLAs to model complex continuous action distributions, thereby achieving more dexterous physical behaviors.
In this work, we further fine-tune the $\pi$-series models with online RL algorithms, enhancing their performance and generalization capabilities through online interaction with the environment.
2.2 Online RL Fine-tuning for VLA Models
Recent research has increasingly focused on enhancing the performance and generalization of VLAs with online RL. For example, SimpleVLA-RL (li2025simplevla), building on the OpenVLA-OFT and GRPO, demonstrated that RL can improve long-horizon planning of VLA models under data scarcity. RL4VLA (rl4vla) empirically evaluated PPO, GRPO, and direct preference optimization (DPO) (rafailov2023dpo) with stage-based sparse rewards, finding PPO to yield superior performance. VLA-RL (lu2025vlarl) proposed a specialized robotic process reward model and enhanced the data processing pipeline. iRe-VLA (guo2025improving) proposed a framework that iterates between RL exploration and SFT updates. RIPT-VLA (tan2025riptvla) applied the REINFORCE leave-one-out (RLOO) (kool2018rloo) algorithm to the QueST (mete2024quest) and OpenVLA-OFT architectures. RLinf-VLA (yu2025rlinf; zang2025rlinf) provides a unified and efficient framework for scalable RL training of VLA models, supporting diverse VLA architectures such as OpenVLA and OpenVLA-OFT, multiple RL algorithms like PPO and GRPO, and various simulators including LIBERO and ManiSkill. These works demonstrate the effectiveness of RL fine-tuning VLA models.
While these approaches demonstrate the potential of applying online RL to VLAs, their application to flow-based VLAs is hindered by the challenge of exact log-likelihood estimation.
2.3 RL Fine-tuning for Flow Models
Integrating RL with flow models is a promising way to transcend the limitations of imitation learning. To this end, Flow-GRPO (liu2025flowgrpo) converts the deterministic ODE into an equivalent SDE to enable stochasticity exploration, a foundation upon which subsequent works like Mix-GRPO (li2025mixgrpo) and TempFlow-GRPO (he2025tempflow) further accelerate training through hybrid ODE-SDE rollouts. ReinFlow (zhang2025reinflow) injects learnable noise into the flow path and transforms it into a discrete-time Markov process with a tractable likelihood for stable policy gradient updates. Flow policy optimization (FPO) (mcallister2025fpo) reframes policy optimization as maximizing the advantage-weighted ratio of the conditional flow matching loss. Policy-Agnostic RL (PA-RL) parl effectively fine-tunes diverse diffusion and Transformer architectures by distilling critic-optimized actions into the policy via supervised learning. Diffusion steering via reinforcement learning (DSRL) (wagenmaker2025steering) refines the flow policy by performing RL in its latent-noise space, rather than modifying the policy parameters themselves.
While prior work has mostly focused on non-robotic tasks or small-scale, single-task robotics, we address the more challenging problem of fine-tuning large-scale flow-based VLAs for complex, multi-task robotic scenarios.
3 Preliminary
3.1 Problem Formulation
We formulate the task as an MDP, defined by a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},P_{0},P_{\text{ENV}},R_{\text{ENV}},\gamma)$. The state $s_{t}\in\mathcal{S}$ is defined as the robot observation $\mathbf{o}_{t}$ and $P_{0}$ denotes the initial state distribution. Given the state, the flow policy predicts an action $a_{t}\sim\pi(\cdot|s_{t})\in\mathcal{A}$, resulting in the state transition $s_{t+1}\sim P_{\text{ENV}}(\cdot|s_{t},a_{t})$ and a reward $R_{\text{ENV}}(s_{t},a_{t})$. The objective is to learn a policy $\pi_{\theta}$ that maximizes the expected $\gamma$-discounted return over a horizon of $T+1$:
$\mathcal{J}(\pi_{\theta})=\mathbb{E}_{\pi_{\theta},P_{0}}\left[\sum_{t=0}^{T}\gamma^{t}R_{\text{ENV}}(s_{t},a_{t})\right].$ (1)
With the policy gradient surrogate (williams1992simple), the gradient of the return expectation can be approximated from sampled trajectories:
$\nabla_{\theta}\mathcal{J}(\pi_{\theta})=\mathbb{E}_{\pi_{\theta},P_{0}}\left[\sum_{t=0}^{T}\nabla_{\theta}\log\pi_{\theta}(a_{t}|s_{t})A(s_{t},a_{t})\right].$ (2)
The advantage function, $A(s_{t},a_{t})=Q(s_{t},a_{t})-V(s_{t})$, measures the relative merit of the action value $Q(s_{t},a_{t})$ over the state value $V(s_{t})$, providing a low-variance signal for the policy update.
3.2 Flow-based Vision-Language-Action Model
A flow-based VLA model $\pi_{\theta}$ is designed to map the observation $\mathbf{o}_{t}$ comprising RGB images, language tokens, and robot proprioception to a sequence of $H$ future actions $\mathbf{A}_{t}=[a_{t,0},...,a_{t,H-1}]$, formulated as $p(\mathbf{A}_{t}|\mathbf{o}_{t})$. Within the model, the VLM extracts features from the visual and language inputs, while the flow matching expert is tasked with generating the actions. Specifically, the model learns a conditional vector field $\mathbf{v}_{\theta}$ that transforms a standard Gaussian noise distribution into the target action $\mathbf{A}_{t}$. This is achieved by minimizing the Conditional Flow Matching (CFM) loss, which aligns the predicted vector field $\mathbf{v}_{\theta}$ with the ground-truth vector field $\mathbf{u}$:
$\mathcal{L}_{\text{CFM}}=\mathbb{E}_{\tau,p(\mathbf{A}_{t},\mathbf{o}_{t}),q(\mathbf{A}_{t}^{\tau}|\mathbf{A}_{t})}\left[\left\|\mathbf{v}_{\theta}(\mathbf{A}_{t}^{\tau},\mathbf{o}_{t})-\mathbf{u}(\mathbf{A}_{t}^{\tau}|\mathbf{A}_{t})\right\|_{2}^{2}\right].$ (3)
[…]
During the inference, the action sequence is generated by first sampling a noise vector $\mathbf{A}_{t}^{0}\sim\mathcal{N}(0,I)$, which is further iteratively refined by integrating the learned vector field $\mathbf{v}_{\theta}$ over a fixed number of steps based on the forward Euler method: $\mathbf{A}_{t}^{\tau+\delta}=\mathbf{A}_{t}^{\tau}+\mathbf{v}_{\theta}(\mathbf{A}_{t}^{\tau},\mathbf{o}_{t})\cdot\delta$.
Figure 2: Two optimization methods in $\pi_{\texttt{RL}}$. Flow-Noise adds learnable noise in a one-layer MDP (Fig.˜3), using the denoised joint likelihood for policy gradient. Flow-SDE builds a two-layer MDP with ODE-to-SDE conversion, and computes the likelihood directly.
4 Methodology
Existing VLA-RL approaches leverage base models such as OpenVLA for discrete actions and OpenVLA-OFT for continuous actions. To compute the action log-likelihood $\log\pi_{\theta}(a_{t}|s_{t})$, discrete models (rl4vla) apply softmax to the output logits, while continuous models (li2025simplevla) treat the action as a Gaussian distribution, employing a prediction head to estimate the variance. As for the flow-based VLAs, directly computing the exact likelihood (hutchinson1989stochastic) is inaccurate with few denoising steps. Moreover, the deterministic nature of its ODE sampling process precludes exploration, making its implementation within RL non-trivial. To this end, we propose Flow-Noise and Flow-SDE, two technical approaches that make flow-based VLAs amenable to RL, as depicted in Fig.˜2.
4.1 Flow-Noise
Inspired by Reinflow (zhang2025reinflow), we incorporate a learnable noise network into the flow matching denoising process and solve the problem within the standard one-layer MDP framework detailed in Sec.˜3.1. By modeling the denoising stage as a discrete MDP, we can directly compute the log-likelihood of the denoised sequence, enabling equivalent policy optimization via RL.
4.1.1 Stochasticity Injection
In Flow-Noise, we parameterize the noise schedule with a neural network, allowing the magnitude of the injected noise to be learned dynamically during training for greater flexibility, as shown in Fig.˜3. We focus on the generation process within a single environment timestep $t$. For notational simplicity, we omit the time subscript $t$, e.g., writing $\mathbf{A}^{\tau}$, and denote the predicted velocity $\mathbf{v}_{\theta}(\mathbf{A}^{\tau},\mathbf{o})$ as $\mathbf{v}^{\tau}$.
The step transition during the denoising process is modeled as an isotropic Gaussian distribution $p(\mathbf{A}^{\tau+\delta}|\mathbf{A}^{\tau})\sim\mathcal{N}(\mu_{\tau},\Sigma_{\tau})$, where the mean is determined by the forward Euler update of the original ODE and the variance is controlled by the learnable noise network $\theta^{\prime}$:
$\begin{cases}\mu_{\tau}=\mathbf{A}^{\tau}+\mathbf{v}^{\tau}\cdot\delta\\ \Sigma_{\tau}=\text{diag}(\sigma_{\theta^{\prime}}^{2})\end{cases}.$ (4)
Here, $\sigma_{\theta^{\prime}}(\cdot)$ is the standard deviation learned from the noise injection network, conditioned on the action $\mathbf{A}^{\tau}$, and the observation $\mathbf{o}$. The noise network is trained jointly with the velocity but discarded after fine-tuning, leaving a deterministic policy for inference.
Figure 3: Illustration for the noise injection on the flow matching, exemplified by $\pi_{0.5}$, which integrates image, language, and state information for unified VLM input.
4.1.2 Log-Likelihood Estimation
The primary challenge in applying policy gradient methods to flow-based VLAs stems from the intractable log-likelihood of the final executed action. In Flow-Noise, we address it by substituting the gradient of the joint log-likelihood of the entire denoising process into the policy optimization objective in Eq.˜2, which is theoretically grounded in Reinflow (zhang2025reinflow).
The inference process for action generation is discretized into $K$ uniform steps, which defines a sequence of time points $\{\tau_{0},\tau_{1},\dots,\tau_{K}\}$. With the step interval defined as $\delta=1/K$, the discrete timestep at the $k$-th point is $\tau_{k}=k\cdot\delta$, starting from $\tau_{0}=0$ and culminating at $\tau_{K}=1$. Given the observation $\mathbf{o}$, the exact and tractable log probability for the entire denoising sequence $\mathcal{A}=(\mathbf{A}^{0},\dots,\mathbf{A}^{1})$ is depicted in Fig.˜2 and formulated as:
$\log\pi(\mathcal{A}|\mathbf{o})=\log\left(\pi(\mathbf{A}^{0}|\mathbf{o})\prod_{k=0}^{K-1}\pi(\mathbf{A}^{\tau_{k+1}}|\mathbf{A}^{\tau_{k}},\mathbf{o})\right).$ (5)
Building on this, we can treat flow-based policy optimization within a standard MDP framework.
4.2 Flow-SDE
Inspired by Flow-GRPO (liu2025flowgrpo), we enhance stochastic exploration by converting the denoising process from ODE into an SDE formulation. We further construct a two-layer MDP to couple the denoising process with the policy-environment interaction following DPPO (ren2024dppo), while leveraging the hybrid ODE-SDE sampling technique to accelerate the training process.
4.2.1 Stochasticity Injection
In Flow-SDE, we convert the deterministic ODE into an equivalent SDE that preserves the marginal probability density of the generated actions, as shown in Fig.˜3.
The deterministic ODE sampling trajectory of the flow matching, especially the Rectified Flow (liu2022recflow), is described by the forward Euler method:
$d\mathbf{A}^{\tau}=\mathbf{v}^{\tau}d\tau.$ (6)
Building on the connection between the probability flow ODE and SDE (song2020score), we can transform the deterministic ODE in Eq.˜6 into an equivalent SDE, with a drift term that corrects the original velocity and a diffusion term that introduces noise:
$d\mathbf{A}^{\tau}=\underbrace{\left(\mathbf{v}^{\tau}-\frac{1}{2}g^{2}(\tau)\nabla\log q_{\tau}(\mathbf{A}^{\tau})\right)d\tau}_{\text{Drift Term}}+\underbrace{g(\tau)d\mathbf{w}}_{\text{Diffusion Term}},$ (7)
where $g(\tau)$ is a scalar function controlling the noise schedule, $\nabla\log q_{\tau}(\mathbf{A}^{\tau})$ is the score function of the marginal distribution $q_{\tau}$ and $d\mathbf{w}$ denotes a Wiener process.
As established in Flow-GRPO, the score function and the velocity field are critically linked by $\nabla\log q_{\tau}(\mathbf{A}^{\tau})=-\frac{\mathbf{A}^{\tau}}{\tau}-\frac{1-\tau}{\tau}\mathbf{v}^{\tau}$. By substituting the score function with the velocity field in Eq.˜7 and setting the noise schedule $g(\tau)$ to $\sigma_{\tau}=a\sqrt{\frac{\tau}{1-\tau}}$ with $a$ controlling the noise level, we derive the final SDE formulation for the flow-matching sampler:
$d\mathbf{A}^{\tau}=\left[\mathbf{v}^{\tau}+\frac{\sigma_{\tau}^{2}}{2\tau}\left(\mathbf{A}^{\tau}+(1-\tau)\mathbf{v}^{\tau}\right)\right]d\tau+\sigma_{\tau}d\mathbf{w}_{\tau}.$ (8)
Discretizing this SDE reveals that the transition probability $p(\mathbf{A}^{\tau+\delta}|\mathbf{A}^{\tau})\sim\mathcal{N}(\mu_{\tau},\Sigma_{\tau})$ is an isotropic Gaussian distribution, with the mean and variance formulated as:
$\begin{cases}\mu_{\tau}=\mathbf{A}^{\tau}+\left[\mathbf{v}^{\tau}+\frac{\sigma_{\tau}^{2}}{2\tau}\left(\mathbf{A}^{\tau}+(1-\tau)\mathbf{v}^{\tau}\right)\right]\cdot\delta\\ \Sigma_{\tau}=\sigma_{\tau}^{2}\delta\cdot\mathbf{I}\end{cases}.$ (9)
4.2.2 MDP Formulation
While Flow-Noise substitutes the joint log-likelihood of the entire denoising sequence for the likelihood of the final executed action, we couple the denoising process of the flow matching with environmental interaction in Flow-SDE. Specifically, we embed the inner MDP defined during the denoising process into the high-level, outer-loop MDP with the environment $\mathcal{M}_{\text{ENV}}$ in Sec.˜3.1, formulating a two-layer MDP as shown in Fig.˜2, with components defined with respect to the environment time $t$ and denoising time $\tau$.
  • State $\bar{s}_{t}^{\tau}=(\mathbf{o}_{t},\mathbf{A}_{t}^{\tau})$ is the tuple of the observation $\mathbf{o}_{t}$ and the action state $\mathbf{A}_{t}^{\tau}$.
  • Action $\bar{a}_{t}^{\tau}$ is defined as the next sampled denoised action in the inner-loop and the executed action for the outer loop:
    $\bar{a}_{t}^{\tau}=\begin{cases}\mathbf{A}_{t}^{\tau+\delta}&\text{if }\tau<1\\ \mathbf{A}_{t}^{1}&\text{if }\tau=1\end{cases},$ (10)
    where $\mathbf{A}_{t}^{\tau+\delta}=\mu_{\tau}+\sigma_{\tau}\sqrt{\delta}\cdot\bm{\epsilon}$, $\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$ is the randomly sampled noise.
  • Transition $\bar{P}(\bar{s}_{t^{\prime}}^{\tau^{\prime}}|\bar{s}_{t}^{\tau},\bar{a}_{t}^{\tau})$ defines how the state evolves, formulated as:
    $\bar{s}_{t^{\prime}}^{\tau^{\prime}}=\begin{cases}(\mathbf{o}_{t},\bar{a}_{t}^{\tau})&\text{if }\tau<1\\ (\mathbf{o}_{t+1},\mathbf{A}_{t+1}^{0})&\text{if }\tau=1\end{cases}.$ (11)
    For $\tau<1$, the inner loop transition $P_{\text{FLOW}}(\cdot)$ occurs between different denoised action states, where the observation $\mathbf{o}_{t}$ remains fixed and the next action state is set by $\bar{a}_{t}^{\tau}=\mathbf{A}_{t}^{\tau+\delta}$.
    For $\tau=1$, the final action $\bar{a}_{t}^{\tau}=\mathbf{A}_{t}^{1}$ interacts with the outer-loop environment, resulting in a new observation $\mathbf{o}_{t+1}$ according to the environment dynamics $P_{\text{ENV}}(\cdot)$. Concurrently, the action state is reset from a standard normal distribution $\mathbf{A}_{t+1}^{0}\sim\mathcal{N}(0,I)$.
  • Reward $\bar{R}(\bar{s}_{t}^{\tau},\bar{a}_{t}^{\tau})$ is granted only upon completion of the denoising process and interaction with the environment:
    $\bar{R}(\bar{s}_{t}^{\tau},\bar{a}_{t}^{\tau})=\begin{cases}0&\text{if }\tau<1\\ R_{\text{ENV}}(\mathbf{o}_{t},\mathbf{A}_{t}^{1})&\text{if }\tau=1\end{cases}.$ (12)
Within the two-layer MDP framework, the problem of estimating the action log-likelihood $\log\pi(a_{t}|s_{t})$ is transformed into estimating $\log\pi(\bar{a}_{t}^{\tau}|\bar{s}_{t}^{\tau})$, which is straightforward to compute due to the Gaussian nature of the transitions.
4.2.3 Hybrid ODE-SDE Sampling
In the formulated two-layer MDP framework, the effective trajectory length is the product of the environment interaction steps and the number of flow matching denoising steps. While this formulation enables RL training for flow-based VLAs, it significantly extends the MDP horizon compared to non-iterative VLA methods, which substantially increases both the training difficulty and the computational time required for optimization.
To this end, we adopt the mixed ODE-SDE rollout strategy, drawing inspiration from the text-to-image generation methods such as Mix-GRPO (li2025mixgrpo) and TempFlow-GRPO (he2025tempflow). Specifically, during the denoising process, a single step is randomly sampled as a stochastic SDE transition governed by $p(\mathbf{A}^{\tau+\delta}|\mathbf{A}^{\tau})\sim\mathcal{N}(\mu_{\tau},\Sigma_{\tau})$, while the remaining steps follow deterministic ODE transitions defined by the update rule $\mathbf{A}^{\tau+\delta}=\mathbf{A}^{\tau}+\mathbf{v}^{\tau}\cdot\delta$.
Under this formulation, we treat the deterministic ODE transition between states as an environment-level wrapper and revise the state transition function of the previous two-layer MDP. Specifically, at each environment step $t$, a denoising time $\tau_{t}$ is randomly selected for the policy’s stochastic injection. The policy $\pi$ acts on this state $\bar{s}_{t}^{\tau_{t}}=(\mathbf{o}_{t},\mathbf{A}_{t}^{\tau_{t}})$, sampling the action $\mathbf{A}_{t}^{\tau_{t}+\delta}$ according to Eq.˜10. The environment wrappers then execute all subsequent deterministic steps, ultimately transitioning to the next observation $\mathbf{o}_{t+1}$ and the next action state $\bar{s}_{t+1}^{\tau_{t+1}}=(\mathbf{o}_{t+1},\mathbf{A}_{t+1}^{\tau_{t+1}})$ at a newly sampled time $\tau_{t+1}$. During this process, the state input and action output of the policy remain consistent with the previous two-layer MDP formulation, thus ensuring theoretical consistency.
4.3 Policy Optimization
4.3.1 Algorithm
Given the formulated flow policy MDP, our objective is to learn the optimal parameters $\theta^{*}$ for the policy $\pi_{\theta}$ that maximizes the expected discounted return $\mathcal{J}(\pi_{\theta})$. To this end, we apply the widely adopted policy gradient algorithm PPO to optimize the policy.
$\pi$-series models (black2024pi_0; intelligence2025pi05) adopt a chunk-based approach for action generation. Specifically, the policy outputs an entire sequence of $H$ future actions $\mathbf{A}_{t}=[a_{t,0},...,a_{t,H-1}]$ in response to each observation. In this approach, we treat the entire sequence as a single macro-step and define its corresponding reward $R_{t}=\sum_{j=0}^{H-1}r_{t,j}$ as the sum of the per-step rewards $r_{t,j}$, referred to as the chunk-level formulation in RLinf-VLA (zang2025rlinf).
To effectively guide policy updates, PPO employs Generalized Advantage Estimation (GAE) (gae) to compute a low-variance estimate of the advantage, estimated as:
$\hat{A}_{t}=\sum_{k=0}^{T-t}(\gamma\lambda)^{k}\mathcal{T}_{t+k},$ (13)
where the TD-error is $\mathcal{T}_{t}=R_{t}+\gamma V(s_{t+1})-V(s_{t})$. Here, $V(\cdot)$ is the state-value function derived from the critic network, $\gamma$ is the discount factor, and $\lambda$ is the parameter that balances the trade-off between bias and variance in the advantage estimate.
PPO constrains policy updates to a small trust region to prevent large, destabilizing updates, with the objective function:
$\mathcal{J}(\pi_{\theta})=\mathbb{E}_{t}\left[\min\left(\rho_{t}(\theta)\hat{A}_{t},\ \text{clip}(\rho_{t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{t}\right)\right],$ (14)
where the clip function, governed by a hyperparameter $\epsilon$, restricts the ratio $\rho_{t}(\theta)$ to the interval $1-\epsilon,1+\epsilon]$ to ensure training stability.
Here, the probability ratio $\rho_{t}(\theta)$ between the updated and old policies takes the form of either:
$\rho_{t}(\theta)=\frac{\pi_{\theta_{\text{new}}}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})}\quad\text{or}\quad\rho_{t}(\theta)=\frac{\pi_{\theta_{\text{new}}}(\bar{a}_{t}^{\tau}|\bar{s}_{t}^{\tau})}{\pi_{\theta_{\text{old}}}(\bar{a}_{t}^{\tau}|\bar{s}_{t}^{\tau})},$ (15)
for the one-layer and two-layer MDP formulations, respectively.
(a) Critic with the action expert, exemplified by $\pi_{0}$.
(b) Critic with the VLM, exemplified by $\pi_{0.5}$.
Figure 4: Illustration of the two critic placement configurations.
4.3.2 Critic Design
Following VLA-PPO works (zang2025rlinf; rl4vla), we employ a shared actor-critic architecture for memory-efficient value prediction as shown in Fig.˜4. However, the two flow-based VLAs process the proprioceptive state differently: in $\pi_{0}$, the state is fed into the action expert model, whereas in $\pi_{0.5}$, it is merged with prompt embeddings within the VLM.
To this end, for the $\pi_{0.5}$ variant, we attach the critic network directly to the VLM output, providing the value estimate $V_{\text{vlm}}(\mathbf{o}_{t})$ conditioned on the integrated image, language, and state inputs. Conversely, for the $\pi_{0}$ variant, achieving the value prediction is non-trivial due to the coupled input structure, where the action expert requires both the noisy action $\mathbf{A}_{t}^{\tau}$ and the state. To this end, we approximate $V_{\text{expert}}(\mathbf{o}_{t})$ by averaging the value estimates across the entire denoising trajectory, formulated as:
$V_{\text{expert}}(\mathbf{o}_{t})\approx\mathbb{E}_{\tau\sim U[0,1]}[V_{\text{expert}}(\mathbf{o}_{t},\mathbf{A}_{t}^{\tau})].$ (16)
5 Experimental Results
5.1 Setup
Benchmarks. We perform experiments based on LIBERO (liu2023libero), ManiSkill (tao2024maniskill3) and MetaWorld (mclean2025meta) benchmarks.
  • LIBERO (liu2023libero) is built on a CPU-based simulation platform MuJoCo and assesses knowledge transfer in robotic multi-task and lifelong learning across four manipulation task suites: Spatial, Object, Goal, and Long.
  • ManiSkill serves as a high-fidelity, GPU-parallelized simulation platform. Within ManiSkill, we adopt the SIMPLER benchmark (SIMPLER) as our primary testbed. To further evaluate the generalization capability of $\pi_{\texttt{RL}}$, we follow the setup of RL4VLA (rl4vla) and construct 4,352 pick-and-place task combinations as an extended benchmark.
  • MetaWorld (mclean2025meta) is a multi-task evaluation benchmark built upon the MuJoCo simulator. To evaluate performance on a broad spectrum of manipulation skills beyond pick-and-place, we utilize the MT50 task set as our testbed.
Flow-based VLAs. We conduct experiments based on $\pi_{0}$ and $\pi_{0.5}$. $\pi_{0}$ introduces the flow-matching action expert (300M) built upon a pre-trained PaliGemma (3B) to leverage broad semantic knowledge from internet-scale data. $\pi_{0.5}$ further utilizes co-training across heterogeneous data sources (e.g., multi-robot data, web data, and high-level semantic predictions) for broader generalization. In addition to the $\pi$-series models, we also conduct experiments on GR00T (bjorck2025gr00t) in Appendix appendix˜C, validating that our algorithm is applicable to other flow-based VLAs.
Implementation Details. Given that pre-trained models often struggle to generalize to task-specific benchmarks, we initiate our process with SFT on expert demonstrations. For the SFT stage, we fine-tune the entire 3.3B model following the official setting. In the subsequent RL stage, we freeze the VLM parameters and exclusively fine-tune the 300M action expert model, driven by GPU memory efficiency and the findings from RL4VLA that RL contributes more significantly to action generalization. We build the whole framework upon the RLinf (yu2025rlinf) codebase, where we adopt a shared, co-located GPU allocation strategy that places the environment, rollout model, and actor model on the same GPU and executes them serially.
For the model configurations, we adhere to the official setting provided by openpi (black2024pi_0; intelligence2025pi05). In these settings, $\pi_{0}$ utilizes image, language, and proprioceptive states as input, whereas $\pi_{0.5}$ notably omits state information for the LIBERO benchmark [^2] https://github.com/Physical-Intelligence/openpi/issues/687. Following this precedent, we consistently omit the state input for $\pi_{0.5}$ during both SFT and RL phases on LIBERO and ManiSkill. Our experiments are conducted on 8 NVIDIA H100 80GB GPUs, and detailed training hyperparameters are available in Appendix Tabs.˜7 and 8.
Table 1: Evaluation results on the LIBERO benchmark, evaluated based on the success rate (%).
Model LIBERO
Spatial Object Goal Long Avg. $\Delta$ Avg.
[…]
Octo 78.9 85.7 84.6 51.1 75.1
OpenVLA 84.7 88.4 79.2 53.7 76.5
$\pi_{\text{fast}}$ 96.4 96.8 88.6 60.2 85.5
OpenVLA-OFT 91.6 95.3 90.6 86.5 91.0
$\pi_{0}$ 96.8 98.8 95.8 85.2 94.2
$\pi_{0.5}$ 98.8 98.2 98.0 92.4 96.9
[…]
$\pi_{0}$ SFT 65.3 64.4 49.8 51.2 57.6
Flow-SDE 98.4 99.4 96.2 90.2 96.1 +38.5
Flow-Noise 99.0 99.2 98.2 93.8 97.6 +40.0
[…]
$\pi_{0.5}$ SFT 84.6 95.4 84.6 43.9 77.1
Flow-SDE 99.6 100 98.8 93.0 97.9 +20.8
Flow-Noise 99.6 100 99.6 94.0 98.3 +21.2
5.2 Main Results
5.2.1 LIBERO
SFT Procedure. The LIBERO benchmark comprises four task suites, each consisting of 10 distinct subtasks. To facilitate few-shot SFT on LIBERO, a minimum of 40 expert demonstration trajectories is necessary to ensure a positive success rate for each subtask across four task suites, thereby guaranteeing a positive optimization signal for the subsequent RL phase.
We perform few-shot SFT following the official training configs provided by openpi. For the $\pi_{0}$ model, we utilized a subset of 58 trajectories, sampled from the total of 1,692 trajectories spanning the four task suites in the official LIBERO SFT dataset [^3] https://huggingface.co/datasets/physical-intelligence/libero, to perform SFT, which served as the initial checkpoint [^4] https://huggingface.co/RLinf/RLinf-Pi0-SFT-Spatial-Object-Goal for subsequent RL training on LIBERO-Spatial, LIBERO-Object and LIBERO-Object task suites. Additionally, a larger pool of 208 trajectories was employed for the LIBERO-Long few-shot SFT [^5] https://huggingface.co/RLinf/RLinf-Pi0-SFT-Long due to the long-horizon and more challenging nature of these tasks. For the $\pi_{0.5}$ model, given its better pretrained checkpoint and training config, we only leveraged 40 trajectories for few-shot SFT, providing a unified checkpoint [^6] https://huggingface.co/RLinf/RLinf-Pi05-SFT across task suites.
RL Procedure. In RL, the VLA model receives a multi-modal input state comprising: an agent-view and a wrist-view (both 224 $\times$ 224 RGB images), natural language guidance, the robot end effector pose, and the gripper state. The model outputs an action to interact with the LIBERO environment, which provides a binary reward of 1 for successful task completion and 0 otherwise.
Experiments. We benchmark the performance of $\pi_{\texttt{RL}}$, which fine-tunes the few-shot SFT $\pi_{0}$ and $\pi_{0.5}$ models with our proposed Flow-Noise and Flow-SDE, against several state-of-the-art VLAs trained on the entire LIBERO dataset, including Octo, OpenVLA, OpenVLA-OFT, $\pi_{\text{fast}}$ (pertsch2025fast), $\pi_{0}$, and $\pi_{0.5}$. We conduct experiments on four LIBERO task suites and report performance as the success rate across all 500 initial states (10 sub-tasks $\times$ 50 states each).
Analysis. As detailed in Sec.˜5.1, our proposed two solutions, Flow-Noise and Flow-SDE, not only achieve comparable performance but also establish a new state-of-the-art by significantly boosting the performance of the few-shot $\pi_{0}$ and $\pi_{0.5}$ SFT models.
For the few-shot $\pi_{0}$ model, the SFT baseline performs poorly, with an average success rate of only 57.6%, indicating that the model struggles with limited demonstration data. Our proposed $\pi_{\texttt{RL}}$ substantially boosts performance, with Flow-SDE and Flow-Noise reaching 96.1% and 97.6%, respectively, and surpassing the full-dataset $\pi_{0}$ SFT baseline of 94.2%.
While the $\pi_{0.5}$ few-shot SFT baseline achieves a decent average performance of 77.1%, it struggles with the challenging LIBERO-Long task, scoring only 43.9%. Our proposed $\pi_{\texttt{RL}}$ framework rectifies this deficiency, boosting the LIBERO-Long success rate from 43.9% to 94.0%, constituting a 50.1% improvement. Notably, despite using only a single trajectory for SFT, $\pi_{\texttt{RL}}$ reaches 98.3% final performance, surpassing the 96.9% full-dataset SFT model.
Discussion on two methods. Flow-SDE and Flow-Noise differ primarily in their noise injection strategy and MDP formulation, with experiments indicating that Flow-Noise marginally outperforms Flow-SDE, a result we attribute to two factors:
  • Noise Injection: Flow-Noise employs a noise network for exploration, complemented by a relative entropy bonus for noise magnitude adaptation, which affords the model finer control during convergence, thus achieving better performance.
  • MDP Formulation: Flow-Noise adopts a one-layer MDP formulation where the log-probability of the executed action is derived from the joint log-probability of the denoised sequence. This formulation endows Flow-Noise with higher data utilization efficiency, leading to faster convergence, as demonstrated in Fig.˜8.
Despite this, the performance discrepancy is still marginal (e.g., $1.5\%$ in $\pi_{0}$ and $0.4\%$ in $\pi_{0.5}$). Additionally, Flow-Noise requires recomputing the entire denoising trajectory for log-likelihood computation. Consequently, the update time per RL training step scales with the number of denoising steps, whereas it remains constant for Flow-SDE due to its mixed ODE-SDE rollout strategy.
5.2.2 ManiSkill
SFT Procedure. Since the SFT dataset provided by RL4VLA lacks the state information required for $\pi_{0}$, we re-synthesized trajectories following their setting using the MPLib motion planning suite (Guo_MPlib), with the final 15 additional frames appended to reinforce the concept of completing a motion.
RL Procedure. In RL, the VLA model receives an input comprising a single 480 x 640 RGB third-person view, a short language instruction, and the current joint pose. The model also receives a structured reward signal from the environment: 1.0 for correct object placement and 0.1 for successful attachment of the gripper to the object, mitigating unwanted throwing behaviors.
Experiments. Based on the models $\pi_{0}$ and $\pi_{0.5}$ models, we empirically validate the performance of Flow-SDE and Flow-Noise against SFT baselines in SIMPLER and a pick-and-place generalization benchmark. Following the default settings of $\pi_{0}$ and $\pi_{0.5}$’s official code base, we include proprioception information in the input to the action expert of $\pi_{0}$ and omit the state input to the VLM of $\pi_{0.5}$.
  • In SIMPLER, the experimental setup comprises an 8-DoF WidowX-250S arm evaluated on four standard tasks: (1) Spoon: placing a spoon on a cloth, (2) Carrot: placing a carrot on a plate, (3) Eggplant: placing an eggplant in a basket, and (4) Cube: stacking a cube. For the SFT stage, we employ a curated dataset in which each task is trained with 144 demonstration episodes.
  • In the generalization test, the policy is prompted to pick from 16 different object types and place them onto 17 different receptacles, distributed across 16 unique table scenes, yielding a total of 4,352 unique task combinations. Given the high complexity of this setting, the SFT data set was prepared with 16,384 episodes, a scale substantially larger than that for SIMPLER tasks.
Analysis. As detailed in Tab.˜2 and Tab.˜3, $\pi_{\texttt{RL}}$ achieves substantial performance improvements in both the SIMPLER and generalization environments. In the SIMPLER environment, $\pi_{\texttt{RL}}$ increases the average success rate of the $\pi_{0}$ model from 67.2% to 86.7%, with three tasks (carrot, eggplant, and spoon) exceeding 90% success. In the training environment of the generalization test, which comprises 4352 task compositions, the performance of $\pi_{0}$ increases from 41.6% to 85.7%, while the $\pi_{0.5}$ model improves from 40.1% to 84.8%. These results demonstrate the effectiveness of $\pi_{\texttt{RL}}$ in a photorealistic environment.
Table 2: Evaluation results on the WidowX SIMPLER benchmark for $\pi_{0}$ and $\pi_{0.5}$.
Model SIMPLER
Carrot Eggplant Spoon Cube Avg.
$\pi_{0}$ SFT 82.7 87.5 61.7 37.1 67.2
Flow-Noise 95.7 96.7 91.6 63.0 86.7
$\Delta$ +13.0 +9.2 +29.9 +25.9 +19.5
$\pi_{0.5}$ SFT 70.6 91.9 43.5 31.0 59.2
Flow-Noise 82.0 98.2 82.8 53.3 79.1
$\Delta$ +11.4 +6.3 +39.3 +22.3 +19.9
Table 3: Evaluation results on the Generalization Test of ManiSkill.
Model IND OOD
Vision Semantic Execution Avg.
$\pi_{0}$ SFT 38.4 32.6 8.4 13.2 18.1
Flow-SDE 78.8 61.1 25.4 31.5 39.3
Flow-Noise 77.8 63.4 23.1 24.2 36.9
$\Delta$ +40.4 +30.8 +16.8 +18.3 +21.3
$\pi_{0.5}$ SFT 40.1 40.2 16.6 22.4 26.4
Flow-SDE 90.9 68.0 34.5 45.4 49.3
Flow-Noise 89.7 69.9 35.5 54.9 53.4
$\Delta$ +50.8 +29.7 +18.9 +32.5 +27.1
Generalization Tests. Following RL4VLA, we further evaluate the model’s generalization across three challenging out-of-distribution (OOD) scenarios: (1) Vision, challenging the model with novel backgrounds and textures; (2) Semantics, probing comprehension with unseen objects, varied instructions, and confounding elements like extra objects or receptacles; (3) Execution, assessing robustness against varied initial states, unseen robot poses, and dynamic disturbances, such as moving target object during execution. In the OOD scenarios detailed in Tab.˜3, we observe that the $\pi_{0}$-SFT model demonstrates strong generalization for visual information. This can be attributed to the robust foundation of its VLM, which allows it to better handle visual disturbances.
However, the semantic performance of $\pi_{0}$ drops dramatically. This degradation is less pronounced when switching to the $\pi_{0.5}$ baseline, a benefit likely stemming from the knowledge generalization of the pre-trained $\pi_{0.5}$ model. Regarding action execution, $\pi_{0}$ exhibits a larger performance drop than $\pi_{0.5}$. We hypothesize that this discrepancy arises from the inclusion of joint angle states as input in $\pi_{0}$, leading to severe overfitting in the control task. In contrast, $\pi_{0.5}$ omits these inputs, thereby avoiding the same degree of performance degradation.
Furthermore, while RL yields significant improvements on in-distribution tasks, we observe its gains are limited in OOD scenarios. We attribute this discrepancy to two factors we aim to address in future work. First, the SFT baseline model itself exhibits substantial performance degradation in OOD settings, which inherently caps the generalization potential achievable by the subsequent RL finetuning. Second, freezing the VLM during the RL stage for training efficiency prevents the model from adapting its visual features to the environment, consequently hindering its visual generalization capabilities.
5.2.3 MetaWorld
SFT Procedure. We perform SFT on the $\pi_{0}$ and $\pi_{0.5}$ models using the official dataset [^7] https://huggingface.co/datasets/lerobot/metaworld_mt50, which consists of 2500 trajectories across 50 different manipulation tasks.
RL Procedure. During the RL procedure, the VLA model processes a multi-modal input comprising a 480 $\times$ 480 RGB agent-view image, language guidance, the robot’s end-effector position, and its gripper state. Based on this input, the model outputs an action to interact with the environment, which in turn provides a sparse reward: 1 for successful task completion and 0 otherwise.
Experiments. We benchmark the performance of $\pi_{\texttt{RL}}$ against Diffusion Policy (diffusion_policy), TinyVLA (wen2025tinyvla), and SmolVLA (shukor2025smolvla). For the performance evaluation, we follow the setup from SmolVLA, i.e., classifying 50 tasks into easy, medium, hard, and very hard four categories according to their difficulties.
Analysis. As detailed in Tab.˜4, RL fine-tuning substantially boosts performance. The $\pi_{0}$ and $\pi_{0.5}$ models achieve average success rates of 85.8% and 70.7%, respectively. This marks a significant improvement over their SFT-only counterparts and surpasses the best-performing baseline SmolVLA of 68.2%, confirming that RL can effectively enhance model capabilities across a diverse range of manipulation task types.
Table 4: Evaluation results on the MetaWorld MT50 benchmark.
Methods MetaWorld
Easy Medium Hard Very Hard Avg. $\Delta$ Avg.
Diffusion Policy 23.1 10.7 1.9 6.1 10.5
TinyVLA 77.6 21.5 11.4 15.8 31.6
SmolVLA 87.1 51.8 70.0 64.0 68.2
$\pi_{0}$ SFT 77.9 51.8 53.3 20.0 50.8
Flow-SDE 92.1 74.6 61.7 84.0 78.1 +27.3
Flow-Noise 91.1 81.8 78.3 92.0 85.8 +35.0
$\pi_{0.5}$ SFT 68.2 37.3 41.7 28.0 43.8
Flow-SDE 86.4 55.5 75.0 66.0 70.7 +26.9
Flow-Noise 86.8 58.1 63.3 56.0 66.1 +22.3
Table 5: Comparison of the PPO and GRPO with Flow-SDE on the LIBERO.
Model LIBERO
Spatial Object Goal Long Avg. $\Delta$ Avg.
$\pi_{0}$ SFT 65.3 64.4 49.8 51.2 57.6
+GRPO 97.8 97.8 83.2 81.4 90.0 +32.4
+PPO 98.4 99.4 96.2 90.2 96.0 +38.4
$\pi_{0.5}$ SFT 84.6 95.4 84.6 43.9 77.1
+GRPO 97.4 99.8 91.2 77.6 91.5 +14.4
+PPO 99.6 100 98.8 93.0 97.9 +20.8
(c) Spatial
(d) Object
(e) Goal
(f) Long
Figure 5: Visual comparison of PPO and GRPO with Flow-SDE $\pi_{0}$ on the LIBERO, demonstrating that PPO outperforms GRPO in terms of convergence performance and training speed.
5.3 Ablation Study
Given that Flow-SDE achieves performance comparable to Flow-Noise while offering higher computational efficiency, we conduct ablation studies with the Flow-SDE method on the LIBERO benchmark to investigate the impact of the RL algorithm, critic design, stochasticity injection strategy, MDP formulation, and various hyperparameters.
5.3.1 RL algorithms
Given the significant performance gains from PPO on the LIBERO benchmark, we also investigated the effectiveness of GRPO (shao2024grpo) (see Appendix A for a detailed description), another widely used policy gradient method applied in VLA+RL training. We compare the performance of PPO and GRPO on both the $\pi_{0}$ and $\pi_{0.5}$ models, with results summarized in Tab.˜5.
We further visualize training curves of PPO and GRPO in Fig.˜5, demonstrating that PPO outperforms GRPO in both final convergence performance and training stability across all four LIBERO task suites.
(a) Eval
(b) Value Loss
(c) Explained Variance
Figure 6: Ablation on the critic structure and placement within Flow-SDE $\pi_{0}$ on the LIBERO-Long, indicating that the critic $V_{\text{vlm}}$ attached after the VLM exhibits superior performance. Furthermore, a four-layer MLP demonstrates stronger regression capability than a one-layer MLP in $V_{\text{expert}}$.
5.3.2 Critic Design
Placement. We compare two critic placement strategies, one positioned after the action expert ($V_{\text{expert}}$) and the other after the VLM ($V_{\text{vlm}}$), with $\pi_{0}$ model on the LIBERO-Long task suite. As illustrated in Fig.˜6, both placements yield comparable performance. However, we observe that $V_{\text{vlm}}$ exhibits slightly superior performance, lower value loss, and higher explained variance, despite not receiving the proprioceptive state as input. This advantage can be attributed to a key difference in their input: $V_{\text{vlm}}$ learns a direct mapping from observation to value, while $V_{\text{expert}}$ must contend with optimization challenges arising from coupled state and noisy action inputs.
Nevertheless, to align with the design of the value function, we maintain the $V_{\text{expert}}$ architecture for the $\pi_{0}$, ensuring that state information is incorporated to calculate the value.
Structure. We investigate a four-layer MLP versus a one-layer MLP, which mirrors the action-projection structure in the action expert. Results in Fig.˜6 indicate that the four-layer MLP leads to a more accurate value approximation, resulting in enhanced performance and training stability.
(a) Train
(b) Eval
Figure 7: Ablation on the injection strategy within Flow-SDE of $\pi_{0}$ on the LIBERO-Long.
5.3.3 Stochasticity Injection
Flow-Noise and Flow-SDE provide two distinct approaches for injecting stochasticity. Specifically, Flow-Noise employs a learnable noise network, while Flow-SDE uses a fixed noise level strategy as illustrated in Fig.˜3. To isolate the impact of the injection strategy, we evaluate these two strategies on the LIBERO-Long task suite, with the same Flow-SDE MDP formulation. Since the fixed noise approach does not incorporate an entropy coefficient, we set the entropy bonus for learned noise to 0 to ensure a fair comparison.
We set the fixed noise level to $a=0.5$, and the lower and upper bounds for the learnable noise log-variance to 0.08 and 0.16, respectively. As depicted in Fig.˜7, two noise strategies exhibit similar train and eval performance at step 0, which indicates comparable noise magnitudes. Furthermore, the converged performance affirms the efficiency of both injection methods.
(a) Eval
(b) Update Time
(c) Explained Variance
Figure 8: Ablation on the MDP formulation within Flow-SDE of $\pi_{0}$ on the LIBERO-Long.
5.3.4 Flow Policy MDP
Flow-Noise and Flow-SDE also differ in their MDP formulation, as shown in Fig.˜2. Built on the standard one-layer MDP, Flow-Noise directly calculates the log-likelihood of the denoised sequence for the policy update. In contrast, Flow-SDE constructs a two-layer MDP by integrating the denoising process with the environment, and further employs a hybrid ODE-SDE sampling technique for acceleration. With the same Flow-SDE noise injection strategy, we evaluate these different frameworks on the LIBERO-Long task suite, as illustrated in Fig.˜8.
While the one-layer formulation converges fastest, all three frameworks achieve similar final performance. In terms of computational cost, the hybrid two-layer paradigm reduces training time by half compared to the standard two-layer approach, thanks to a shorter effective MDP chain that lowers the computational cost per RL update. Moreover, we observe that the one-layer MDP shows no significant speed advantage over the standard two-layer model, as its update stage necessitates re-computing the entire denoising trajectory to calculate the log-likelihood, resulting in comparable computational overhead.
5.3.5 Hyper-parameters
Building on the Flow-SDE with $\pi_{0}$, we investigate the influence of the noise level, denoise step, and action chunk on the LIBERO-Spatial benchmark. We denote the train stage as the phase where the policy generates stochastic actions for exploration, whereas the evaluation stage involves generating deterministic actions. The train and eval success rates for the SFT baseline and the RL fine-tuned model after 100 training steps are presented in Tab.˜6.
Table 6: Ablation study of hyperparameters for Flow-SDE on the LIBERO-Spatial. Performance is reported as task success rate (%). “Train” refers to policy performance during the stochastic rollout phase, whereas “Eval” refers to performance during the deterministic evaluation phase.
Models Stage Hyperparameters
Noise Level Denoise Step Action Chunk
0.2 0.5 0.8 1 2 4 8 5 10 20
SFT Train 62.3 56.0 46.6 9.4 28.3 56.1 62.6 56.0 60.7 70.3
Eval 65.2 65.2 65.2 63.8 64.9 65.2 63.2 65.2 70.5 72.6
RL Train 59.5 93.5 95.3 73.8 90.8 93.5 84.3 93.5 93.3 87.5
Eval 73.1 94.5 98.1 88.5 97.0 94.5 86.7 94.5 95.5 89.2
Noise Level. The noise level $a$ in the Flow-SDE is defined in Eq.˜8, which governs the noise injection magnitude during the denoising process. From Tab.˜6, we observe that the SFT baseline’s eval performance is identical across all noise levels as it relies on deterministic ODE sampling. Its training performance, however, degrades as the noise level increases, which is intuitive as higher noise can disrupt the flow path and lead to an inaccurate marginal action distribution.
Extending this analysis to the RL fine-tuning stage reveals a key trade-off: while lower noise levels mitigate performance degradation induced by policy exploration, the capacity for RL refinement is correspondingly constrained. This trade-off is empirically validated in Fig.˜9, which indicates that training with the minimal noise level $a=0.2$ exhibits instability, manifesting as a significantly higher clip fraction. We attribute this instability to the substantially larger gradient magnitudes induced by the low noise level.
(a) Train
(b) Eval
(c) Clipped Fraction
Figure 9: Ablation on the noise level $a$, conducted with the Flow-SDE $\pi_{0}$ on the LIBERO-Spatial.
Denoise Step. The denoise step $K$ defines the number of discretization steps for action generation and is critical for controlling the fidelity of the ODE-to-SDE transition in Eq.˜8. In Tab.˜6, we observe that while all configurations start with similar eval performance, the train success rate plummets at $K=1$, indicating a significant ODE-to-SDE discretization error.
However, as in our noise-level analysis, a larger $K$ is not necessarily optimal. As shown in Fig.˜10, a larger $K$ introduces a clear trade-off: it yields higher rollout performance but complicates the training process due to an increased number of denoising steps.
(a) Train
(b) Eval
Figure 10: Ablation on the denoise step, conducted with the Flow-SDE $\pi_{0}$ on the LIBERO-Spatial.
(a) Eval
(b) Explained Variance
Figure 11: Ablation on the chunk size, conducted with the Flow-SDE $\pi_{0}$ on the LIBERO-Spatial.
Action chunk. The action chunk refers to the number of consecutive actions the policy executes within a single observation. We ablate the action chunk size across 5, 10, and 20, with results presented in Tab.˜6 and further visualized in Fig.˜11.
While a larger chunk size yields a marginal performance improvement, it also reduces the frequency of policy-environment interactions and hinders accurate reward credit assignment. These factors contribute to less reliable advantage estimation, as reflected in the explained variance metric. Consequently, while a large chunk size may provide a stronger SFT baseline, it ultimately constrains the potential gains from subsequent RL fine-tuning. In conclusion, our analysis reveals a consistent trade-off:
Caveat: Hyperparameters optimized for rollout might induce training instability, impeding potential performance gains from RL.
Therefore, a careful selection of these parameters is essential to achieve a suitable balance between train performance and a stable training process.
5.4 Insights from Large-Scale Training
In this subsection, we elaborate on some empirical insights we gained during RL training.
Hyperparameters. According to the hyperparameters ablation detailed in Sec.˜5.3.5, the performance disparity between the train and eval performance of the initial SFT checkpoint warrants close attention. If this disparity is significant, we recommend either reducing the noise magnitude or increasing the number of denoising steps to mitigate the performance degradation caused by the discrepancy between deterministic and stochastic action generation. Furthermore, as previously established, lower noise levels yield larger gradients, requiring a smaller learning rate to maintain training stability.
We also observed that when train performance improves steadily while eval performance oscillates, increasing the number of denoising steps can help alleviate this, benefiting from reduced divergence in the action distributions between the deterministic and stochastic action generation processes. Regarding the action chunk, we empirically found that long-horizon tasks benefit from larger chunk sizes. For instance, we set the chunk size to 10 for LIBERO-Long and 5 for the other sub-tasks.
Training. In our $\pi_{0.5}$ experiments on the LIBERO-Long benchmark, we observed that the Kullback–Leibler (KL) divergence metric increased steadily throughout training, potentially leading to instability. We mitigated this issue by implementing a learning rate scheduler with cosine annealing. As demonstrated in Fig.˜12, this scheduler effectively prevents the KL divergence from escalating, thereby stabilizing the training process.
(a) Eval
(b) KL Divergence
Figure 12: Ablation study on the learning rate scheduler. The experiment is conducted with Flow-SDE $\pi_{0.5}$ on the LIBERO-Long benchmark, demonstrating that the scheduler alleviates over-optimization and stabilizes the training process.
Critic. In our ManiSkill experiments, we observe that policy evaluation performance exhibits an initial dip before improving for both $\pi_{0}$ and $\pi_{0.5}$ models, as shown in Fig.˜14. We attribute this transient degradation to the critic providing inaccurate signals during its warm-up phase. The subsequent eval improvement correlates directly with the critic’s value estimations stabilizing, as evidenced by the rising explained variance.
Temporal Efficiency We also study how the rollout of RL in a physical simulator helps shape the policy to achieve expert-level temporal efficiency. We analyze the expert motion planning data used for SFT and then tracked the average episodes lengths during the RL training of the $\pi_{0.5}$ model using our methods. As shown in Figure Fig.˜13, the SFT-initialized policy exhibits significantly longer episodes due to execution errors. In contrast, $\pi_{0.5}$ achieves episode lengths that converge to the expert range after RL training, demonstrating a substantial improvement in temporal efficiency.
We attribute this convergence to two factors: (1) RL’s error-correction capability, which helps the policy succeed in a more diverse distribution, and (2) our partial reset mechanism with discounted reward, where faster task completion leads to more resets and higher cumulative reward between updates.
Figure 13: Episode length: $\pi_{0.5}$ RL training in multi-object pick-and-place environment
(a) Eval
(b) Explained Variance
Figure 14: Flow-Noise Training curve in ManiSkill generalization test.
5.5 Extension: Fine-tune VLM and Action Expert Simultaneously
In our previous experiments, the VLM is frozen, and the optimization is confined exclusively to the action expert during RL. In this subsection, we aim to investigate the role of the VLM during RL. Specifically, we employ Low-Rank Adaptation (LoRA) (hu2022lora) for the VLM, facilitating its joint optimization with the action expert. We set the LoRA rank to $r=32$ and the scaling parameter to $\alpha=32$, while the action expert remains fully trainable.
We conduct experiments with the $\pi_{0}$ model with Flow-SDE on the LIBERO-Long benchmark, comparing three distinct configurations: 1) VLM frozen baseline ($5\times 10^{-6}$ learning rate, 4 updates per epoch), 2) VLM LoRA-I ($5\times 10^{-6}$ learning rate, 4 updates per epoch), and 3) VLM LoRA-II with conservative update training config ($1\times 10^{-6}$ learning rate, 2 updates per epoch). As presented in Fig.˜15, the VLM LoRA-II configuration achieves a learning trajectory comparable to the VLM frozen baseline. This empirical observation yields two critical inferences: First, the benefit of fine-tuning the VLM on the LIBERO benchmark is not evident; Second, fine-tuning VLM together with the action expert requires a more conservative optimization configuration for training stability. We conjecture the limited performance gain attributable to the limited scene variability within LIBERO, for which the pretrained VLM representations are already sufficiently robust.
(a) Eval
(b) KL Divergence
Figure 15: Ablation study on VLM Effectiveness during RL. The experiment is conducted with Flow-SDE $\pi_{0}$ on the LIBERO-Long benchmark. We compare the performance of a frozen VLM baseline (learning rate $5\times 10^{-6}$, 4 updates per epoch) against two LoRA-tuned VLM configurations: LoRA-I (using the same training config) and LoRA-II (a more conservative setting with learning rate $1\times 10^{-6}$ and 2 updates per epoch).
6 Conclusion
We introduce $\pi_{\texttt{RL}}$, the first framework that enables flow-based VLAs, $\pi_{0}$ and $\pi_{0.5}$, to be fine-tuned with PPO. We tackle the fundamental challenge of intractable log-likelihoods in flow matching and propose two technical solutions, Flow-Noise and Flow-SDE, which differ in their stochasticity injection strategies and MDP formulations. Our extensive experiments on the challenging LIBERO and ManiSkill benchmarks demonstrated that $\pi_{\texttt{RL}}$ achieves significant performance improvements over SFT baselines.
7 Limitations and Future Work
Noise Injection. Our current noise injection strategy exhibits some train performance drop during the ODE-to-SDE conversion. Flow-CPS (wang2025coefficients) attributes this loss to numerical error and proposes an improved coefficients-preserving sampling method. In our experiments, we attempted this configuration. Consistent with our hyperparameter ablation, our experiments showed that while this configuration mitigated the ODE-SDE precision error, it yielded limited RL improvement. Nevertheless, we argue that improving the noise injection strategy holds significant potential, specifically converting the ODE formulation to an SDE formulation while preserving the action distribution undisturbed.
Training Acceleration. Our current implementation of the mixed ODE-SDE rollout is simplistic in Flow-SDE, i.e., it randomly selects one denoising step as an SDE step, while all other steps remain ODE steps. We posit that future investigations into mixed ODE-SDE rollouts, leveraging advances in accelerating flow-based image generation (li2025mixgrpo; he2025tempflow; liu2025flowgrpo; li2025branchgrpo), could further enhance Flow-SDE, leading to faster training and improved performance.
Generalization. Our experiments in the Maniskill OOD tests indicate that the semantic generalization capabilities of the SFT and RL models remain limited. We aim to investigate and improve this issue in future studies.
Real-world Experiment. Our current experiments are evaluated solely in simulated environments, lacking empirical validation in a physical system. We plan to extend this research by applying our RL methodology to real-world tasks in the future.
Appendix A Algorithm Details
GRPO is a critic-free method that estimates the advantage by normalizing rewards within a group of rollouts from the same state. In our robotics MDP task, for each initial state, we use the policy $\pi_{\theta}$ to sample a group of $G$ trajectories, resulting in $G$ sparse terminal rewards $\{\mathcal{R}^{(j)}\}_{j=1}^{G}$ denoting the binary success of the task. The advantage for the $i$-th trajectory, $\hat{A}^{(i)}$, is then calculated based on the group-wise reward normalization:
$\hat{A}^{(i)}=\frac{\mathcal{R}^{(i)}-\text{mean}(\{\mathcal{R}^{(j)}\}_{j=1}^{G})}{\text{std}(\{\mathcal{R}^{(j)}\}_{j=1}^{G})}$ (17)
where $\mathcal{R}^{(i)}$ is the terminal reward for the $i$-th trajectory, and the mean and standard deviation are computed over the group of $G$ trajectories. Since the reward is only granted at the end of an episode, the advantage estimate remains constant across all timesteps within that trajectory.
Appendix B Experiment Details
We record the training hyperparameters used to train both $\pi_{0}$ and $\pi_{0.5}$ on each LIBERO task, and present them in Tabs.˜7 and 8.
Table 7: Hyperparameters of Flow-Noise and Flow-SDE with PPO across LIBERO tasks.
Parameters Algorithms and tasks
$\bm{\pi_{0}}$ $\bm{\pi_{0.5}}$
Spatial Object Goal Long Spatial Object Goal Long
Train epochs 400 400 400 400 400 400 400 400
Batch size 2048 2048 2048 2048 2048 2048 2048 2048
Update epochs 2 2 4 4 1 1 3 4
Actor lr 1e-5 5e-6 5e-6 5e-6 5e-6 5e-6 5e-6 5e-6
Critic lr 1e-4 1e-4 1e-4 1e-4 1e-4 1e-4 1e-4 1e-4
Scheduler False False False False False False False True
Reward discount rate $\gamma$ 0.99 0.99 0.99 0.99 0.99 0.99 0.99 0.99
GAE $\lambda$ 0.95 0.95 0.95 0.95 0.95 0.95 0.95 0.95
Clip ratio $\epsilon$ 0.2 0.2 0.2 0.2 0.2 0.2 0.2 0.2
Interaction steps 240 320 320 480 240 320 320 480
Parallel environments 64 64 64 64 64 64 64 64
Rollout epochs 8 8 8 8 8 8 8 8
Action chunk $H$ 5 5 5 10 5 5 5 10
Denoise steps 4 4 4 4 3 5 5 5
Noise level $\sigma$ (Flow-SDE) 0.5 0.5 0.5 0.5 0.5 0.3 0.3 0.5
Max log-var (Flow-Noise) 0.16 0.16 0.16 0.16 0.10 0.10 0.10 0.10
Min log-var (Flow-Noise) 0.08 0.08 0.08 0.08 0.04 0.04 0.04 0.04
Entropy bonus (Flow-Noise) 0.005 0.005 0.005 0.005 0.005 0.005 0.005 0.005
Table 8: Flow-Noise and Flow-SDE Hyperparameters of in ManiSkill tasks.
Parameters Algorithms and tasks
$\bm{\pi_{0}}$ $\bm{\pi_{0.5}}$
Eggplant Carrot Spoon Cube Generalization Eggplant Carrot Spoon Cube Generalization
SFT train steps 1000 100040 1000 1000 1000 1000 1000 1000 1000 1000
RL train steps 40 40 40 130 150 40 40 40 70 150
Global Batch size 2560 2560 2560 2560 5120 2560 2560 2560 2560 5120
Update epochs 4 4 4 4 4 4 4 4 4 5
Actor lr 5.6e-6 5.6e-6 5.6e-6 5.6e-6 7.91e-6 5.6e-6 5.6e-6 5.6e-6 5.6e-6 7.91e-6
Critic lr 1.1e-4 1.1e-4 1.1e-4 1.1e-4 1.55e-4 1.1e-4 1.1e-4 1.1e-4 1.1e-4 1.55e-4
Scheduler False False False False False False False False False False
Reward discount rate $\gamma$ 0.99 0.99 0.99 0.99 0.99 0.99 0.99 0.99 0.99 0.99
GAE $\lambda$ 0.95 0.95 0.95 0.95 0.95 0.95 0.95 0.95 0.95 0.95
Clip ratio $\epsilon$ 0.2 0.2 0.2 0.2 0.2 0.2 0.2 0.2 0.2 0.2
Interaction steps 48 48 48 48 48 48 48 48 48 48
Parallel environments 256 256 256 256 320 256 256 256 256 320
Rollout epochs 1 1 1 1 1 1 1 1 1 1
Action prediction horizon $H$ 8 8 8 8 8 8 8 8 8 8
Action replan horizon $H^{\prime}$ 5 5 5 5 5 5 5 5 5 5
Denoise steps 4 4 4 4 4 4 4 4 4 4
Noise level $\sigma$ (Flow-SDE) 0.5 0.5 0.5 0.5 0.5 0.5 0.5 0.5 0.5 0.5
Max log-var (Flow-Noise) 0.16 0.16 0.16 0.16 0.16 0.10 0.10 0.10 0.10 0.10
Min log-var (Flow-Noise) 0.08 0.08 0.08 0.08 0.08 0.04 0.04 0.04 0.04 0.04
Entropy bonus (Flow-Noise) 0.005 0.005 0.005 0.005 0.005 0.005 0.005 0.005 0.005 0.005
w
Table 9: $\pi_{0}$, $\pi_{0.5}$ Generalization Test
Environment Variation-Version-Type $\pi_{0}$-SFT $\pi_{0}$-RL $\pi_{0}$-RL $\pi_{0.5}$-SFT $\pi_{0.5}$-RL $\pi_{0.5}$-RL
Flow-SDE Flow-Noise Flow-SDE Flow-Noise
In distribution Main-v3-train 38.42 78.83 77.76 40.06 90.85 89.65
Visual-language
Variations
Instruct-v1-test 30.10 64.58 66.46 46.56 76.98 85.73
VisionImage-v1-test 38.33 68.75 71.67 46.25 78.75 83.13
VisionTexture03-v1-test 35.10 66.04 66.77 36.67 69.58 75.00
VisionTexture05-v1-test 31.04 55.83 60.52 32.71 58.02 62.19
VisionWhole03-v1-test 35.42 62.40 68.96 40.10 69.58 71.56
VisionWhole05-v1-test 28.54 48.96 53.85 30.73 55.00 56.98
Semantic Reasoning
(object/receptacle
confounders)
MultiCarrot-v1-test 7.81 28.23 23.02 16.67 36.77 38.23
MultiCarrot-v1-train 12.50 36.46 31.77 28.23 49.48 50.10
MultiPlate-v1-test 5.00 16.35 18.33 11.77 29.38 28.33
MultiPlate-v1-train 7.29 20.52 19.58 9.69 22.29 25.42
Action Execution PositionChangeTo-v1-test 9.58 17.40 10.94 13.54 36.25 54.69
Position-v1-test 16.88 45.63 37.50 31.15 54.48 55.00
Appendix C Additional Results: RL for GR00T N1.5
C.1 Setup
GR00T N1.5. We conduct additional experiments based on the GR00T N1.5 model. GR00T N1.5 is an open-source foundation model  [^8] https://github.com/NVIDIA/Isaac-GR00T tailored for generalist humanoid robot reasoning and manipulation. Designed to enable cross-embodiment adaptability, the model accepts multimodal inputs—including natural language instructions and visual observations—and integrates robot state information (e.g., joint positions, end-effector poses) to generate continuous motor actions for diverse tasks and environments. Its neural architecture combines a vision-language model (Eagle 2.5) optimized for grounding and physical understanding with a diffusion transformer (DiT) head for action denoising. It supports multiple robot embodiments via specialized heads—including humanoid robots with dexterous hands, single-arm robots, and humanoid robots with grippers. For the critic placement, we estimate the value across the entire denoising trajectory, attaching the critic network to the action head. The figure illustration is shown in Fig.˜16.
Benchmark. We use LIBERO as the evaluation benchmark. Evaluates the model performance across four manipulation task suites: Spatial, Object, Goal, and Long.
Implementation Details. Similar to the $\pi_{0}$ implementation, we initiate our process with SFT on expert demonstrations. For the SFT stage, we fine-tune the entire model following the official setting. In the subsequent RL stage, we exclusively fine-tune the action expert model while keeping the vision-language model parameters fixed. A critical methodological detail during the RL stage is the replacement of the original dropout layers in the expert model with identity layers. Dropout layers are known to destabilize online RL training. Specifically, they introduce an unintended change in the effective policy, transforming the stable probability ratio update from:ρ_t(θ) = πθnew(at|st)πθold(at|st)to the unstable form:ρ_t(θ) = παnew(at|st)πθold(at|st),where $\rho_{t}(\theta)$ is the probability ratio, and $\alpha_{\text{new}}$ signifies the policy after the update and the non-deterministic effect of the dropout mask. This stochasticity, which effectively changes the model’s structure in addition to the per-step policy update, significantly deteriorates training stability. For the model configurations, we adhere to the official settings. Our experiments are conducted on 8 NVIDIA H100 80GB GPUs, and detailed training hyperparameters are exactly the same as $\pi_{0}$, listed in Tab.˜7. The RL training method selected is Flow-SDE.
Figure 16: Illustration for the architecture of GR00T-N1.5.
Table 10: Comparison of the PPO and GRPO with Flow-SDE on the LIBERO.
Model LIBERO
Spatial Object Goal Long Avg. $\Delta$ Avg.
$GR00T$ SFT 41.4 58.6 48.2 61.9 52.5
+PPO 92.5 96.2 84.3 86.6 89.9 +37.4
C.2 Results
This evaluation adopts all the hyperparameter settings of $\pi_{0}$ without further tuning to show the broad applicability and robustness of the proposed RL training method. The evaluation results are shown in Tab.˜10 For the few-shot model, the SFT baseline performs poorly, with an average success rate of only 52.5%, indicating that the model struggles with limited demonstration data. Our proposed RL training method substantially boosts performance, reaching 89.9% with Flow-SDE.
The results presented above utilize the identical hyperparameter settings as $\pi_{0}$. These findings primarily serve to demonstrate the broad applicability and inherent robustness of the proposed RL training methodology. Further optimization through parameter tuning is likely to yield enhanced model performance.